Consider the following system of linear equations:

-x1+2x2+x3=4

5x1-2x2+3x3=-28

2x1-x2+4x3=-23


Use Cramer’s rule to solve for x^2. (When working out determinants, indicate

with which row or column co-factor expansion is done.)


1
Expert's answer
2021-03-31T10:48:20-0400

The determinant of the coefficient omitting the constant,

"\begin{vmatrix}
 -1 & 2 & 1\\
 5 & -2 & 3\\
 2 & -1 & 4
\end{vmatrix}\\"

Using the row 1 co-factor expansion

"D_0=-1\begin{vmatrix}
 -2 & -3 \\
 -1 & 4
\end{vmatrix}-2\begin{vmatrix}
 5 & 3 \\
 2 & 4
\end{vmatrix}+1\begin{vmatrix}
 5 & -2 \\
 2 & -1
\end{vmatrix}\\
D_0=
-1(-11)-2(14)+1(-3)\\
D_0=-20\\
\text{the determinant of the coefficient omitting},x^2 term\\
D_2=
\begin{vmatrix}
 -1 & 1 & -4\\
 5 & 3 & 28\\
 2 & 4 & 23
\end{vmatrix}\\
\text{Row 1 expansion}
D_2=-1\begin{vmatrix}
 3 & 28 \\
 3 & 23
\end{vmatrix}-1\begin{vmatrix}
 5 & 28\\
 4 & 23
\end{vmatrix}-4\begin{vmatrix}
 5 & 3\\
 2 & 4
\end{vmatrix}\\
D_2=
-1(-43)-1(3)-4(14)\\
D_2=-16\\
x^2=\frac{D_2}{D_0}\\
x^2=\frac{-16}{-20}\\
x^2=\frac{4}{5}"